### v1

# Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions

Abstract
Emerging research has highlighted that artificial intelligence based multimodal fusion of digital pathology and transcriptomic features can improve cancer diagnosis (grading/subtyping) and prognosis (survival risk) prediction. However, such direct fusion for joint decision is impractical in real clinical settings, where histopathology is still the gold standard for diagnosis and transcriptomic tests are rarely requested, at least in the public healthcare system. With our novel diffusion based crossmodal generative AI model PathoGen, we show that genomic expressions synthesized from digital histopathology jointly predicts cancer grading and patient survival risk with high accuracy (state-of-the-art performance), certainty (through conformal coverage guarantee) and interpretability (through distributed attention maps). PathoGen code is available through GitHub at https://github.com/Samiran-Dey/PathoGen for open use by the research community.
1 Introduction
Histopathological assessment is the gold standard for cancer diagnosis and provides information essential for accurate staging to inform prognosis. However, aging populations, strained healthcare services and the increasing diagnostic and prognostic complexity of modern medicine are contributing to an unsustainable workload for practising histopathologists Kelly et al. [2020]. Modern artificial intelligence (AI) algorithms, particularly deep learning models, can predict cancer grade and sub-types from digital hematoxylin and eosin (H&E) stained pathology slides, complementing and supporting pathology workflows and potentially reducing the high workload burden Couture et al. [2018], Arvaniti et al. [2018], Elazab et al. [2024], Zanjani et al. [2018], Das et al. [2022], Wang et al. [2016]. Very recently the emergence of pathology foundation models using large vision transformers (ViT) has greatly increased model flexibility, enabling support with a wide variety of downstream tasks and providing greater utility for pathologists Chen et al. [2024], Lu et al. [2023, 2024].
Figure 1: Data and methodology overview. (a-b) (i) Distribution of data across demographic categories (ii) Distribution of data across train - validation - calibration - test splits (iii) Number of genes in different gene groups. (c) Pipeline for the proposed methodology. Our novel diffusion-based crossmodal generative model, PathoGen, synthesizes transcriptomic features from whole slide image patch embeddings obtained using a state-of-the-art foundation model from the whole slide images. Multimodal prediction is performed for both diagnosis (cancer grading) and prognosis (survival risk) using the synthesized transcriptomic data and histopathology images to predict the added value of transcriptomic data, for each patient. Uncertainty quantification provides a patient-level estimate of model reliability and provides clinicians with survival risk bounds and tumour grade prediction sets which are guaranteed to contain the ground truth with a specified probability. Distributed predictions also provide a window to understand the intra-tumour heterogeneity.
Classically histopathology has relied on microscopic assessment of H&\&&E stained tissue sections, often complemented by targeted panels of immunohistochemical and other special stains, to confirm and refine diagnoses. As technology has advanced, a broader range of molecular tests have become routine for assessing certain tumours. These include genetic analyses such as IDH mutation status, which is required to appropriately subtype gliomas according to the WHO Reuss [2023] and EGFR mutation status which is essential to guide therapy decisions in non-small cell lung cancers Wang and Fang [2014]. More recently, a range of transcriptomic tests have been developed, which show clear evidence for guiding treatment decisions in specific malignancies Sparano et al. [2018]. Following extensive clinical evaluation, several of these tests have been granted regulatory approval for clinical use Tsakiroglou et al. [2023]. However, guidelines, especially in public healthcare systems, strictly limit the application of these transcriptomic tests to a small minority of patients, largely due to their high financial cost and infrastructural requirements Harnan et al. [2019].
Recently, AI models employing a multimodal fusion of digital pathology and transcriptomic features, trained on large public datasets, like The Cancer Genomic Atlas (TCGA), have led to significant improvements in the predictive accuracy of both grading and survival risk estimation, compared to unimodal models employing digital H&\&&E slides alone Mobadersany et al. [2018a], Chen et al. [2020, 2021]. Though these powerful models have the potential to convey significant health benefits, they are developed in a setting disjoint from the current reality of the clinic. Translation of these models into most healthcare systems is infeasible at present, as the required transcriptomic data is very seldom collected. Future translation requires rigorous, costly and widespread clinical evaluation, to determine which patients would benefit from these models sufficiently to justify the cost of transcriptomic data.
Crossmodal generative AI models can be trained to synthesize one unavailable data modality (transcriptomics) from routinely available modalities (digital pathology). Coupled with a multimodal AI model, this could facilitate a cost-effective screening tool to estimate the added value of collecting transcriptomic data for diagnostic/prognostic assessment of each individual patient. Such an approach could support cost-efficient clinical evaluation, and sustainable translation of multimodal AI models to support pathology workflows.
Here we introduce a novel, diffusion based crossmodal generative model, PathoGen, that synthesizes transcriptomic data from digital images of H&\&&E stained slides. We develop our model on two publicly available multimodal data cohorts (glioma Network [2008, 2015] and renal cancer Network [2013]) from TCGA. We find that generated transcriptomic data are highly similar to the corresponding real transcriptomic data, in a hold-out validation set. We further demonstrate that the predictive performance of both diagnosis (cancer grading) and prognosis (survival risk) significantly improves when the synthesised transcriptomic data is combined with features learned from the corresponding digital pathology image, using a state-of-the-art ViT-based model through a co-attention mechanism.
For practical utility, our model must be trusted by clinicians and attune to emerging regulation Commission [2021]. We thus provide transparency via detailed attention based heatmaps for both grading and survival risk prediction, allowing pathologists to visually affirm the clinical relevance of the regions contributing most towards the AI decision. We provide uncertainty quantification, via conformal analysis Angelopoulos and Bates [2021], to allow pathologist users, a measure of certainty in added value of the synthesized transcriptomic data for each patient. Finally, to evaluate algorithmic fairness, we apply our conformal uncertainty quantification to assess model performance analysis across various patient demographics. Our work builds on preliminary studies in crossmodal prediction Chen et al. [2024, 2020, 2021], to provide a road map for sustainable, safe and cost-effective translation of multimodal AI models supporting cancer diagnostics/prognostics, to the clinic.
Figure 2: Evaluation of synthesized transcriptomic data and explainability. (i) comparison between real and synthesized gene expression levels for different gene groups. (ii) plot of significant performance improvement on using synthesized transcriptomic data and not using them and closeness of using synthesized and real transcriptomics for gradation (AUC) and survival risk estimation (C Index). (iii) comparison of distributed prediction with non-distributed WSI prediction. (iv) comparison of co-attention maps generated using synthesized and real transcriptomic data. (v) percentage contribution of gene groups in co-attention.
2 Results
2.1 Dataset and study design
Our experiments use publicly available data from TCGA  [^1] https://www.cancer.gov/tcga, specifically two independent cancer cohorts - Glioblastoma Multiforme and Brain Lower Grade Glioma (TCGA-GBMLGG) and Kidney Renal Clear Cell Carcinoma (TCGA-KIRC) Network [2013]. For TCGA-GBMLGG, 653 whole slide images (WSIs) coming from 532 cases are used to train, 69 WSIs from 56 cases to validate, 104 WSIs from 82 cases to calibrate and 86 WSIs from 75 cases to test (Figure 1.a.ii) of which 562 cases belong to TCGA-LGG Network [2015] and 183 cases belong to TCGA-GBM Network [2008]. For TCGA-KIRC Network [2013], we use 344 WSIs from 337 cases to train, 42 WSIs from 34 cases to validate, 48 WSIs from 45 cases to calibrate and 51 WSIs from 46 cases to test (Figure 1.b.ii). Cases with missing grade, survival data or transcriptomic data are excluded. The data distribution across various demographic groups like gender, age, censorship status, etc. varies for both cohorts (Figures 1.a.i, 1.b.i).
Transcriptomic features are selected to match those employed in the successful multimodal models Chen et al. [2020, 2021]. As proposed previously Chen et al. [2021], transcriptomic data for each patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors  [^2] https://github.com/mahmoodlab/MCAT/blob/master/datasets_csv_sig/signatures.csv. The number of available gene expression levels for each group are 84, 314, 498, 424, 1396 and 428 respectively for TCGA-GBMLGG resulting in a total of 3144 genes (Figure 1.a.iii). Whereas for TCGA-KIRC, the number of available gene expression levels for each group are 145, 522, 669, 681, 2366, and 990 respectively resulting in a total of 5373 genes (Figure 1.b.iii). It is to be noted that there is some overlap (by design) between the gene sets due to the intersecting functionalities they represent. Since our model processes the gene expression levels of all genes in a gene set together, the gene expression embedding corresponding to a gene group is distinct from the embeddings of the other gene groups even if the intersection of the gene sets is not null. Hence, in our paper, we comment on the importance of different gene groups to explain the grade and survival risk predictions and not on a particular gene. When discussing gene expression levels throughout this paper, we refer to transcript abundance, rather than protein levels.
Our methodology synthesizes transcriptomic data from WSIs using our diffusion based generative model, PathoGen. Synthesized transcriptomes are then used for automated cancer gradation and survival risk estimation to assess whether transcriptomic tests should be requested (Figure 1.c). We use transcriptomic data (both real and synthesized) at two stages in the predictive model (MCAT_GR) - to co-attend the WSI patch embeddings and learn a combined feature map that is suitable to predict both grade and survival and to predict the survival risk of a patient in a learnable ensemble method along with the co-attended feature maps. We perform extensive experiments using synthesized transcriptomic data to decide on this model setting (supplementary Table 1). The model is trained end-to-end for performing both gradation and survival risk estimation together using the loss function, $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$, where $L_{survival}$ is the loss for survival risk estimation, $L_{grade}$ is the loss for gradation and $\lambda$ is the loss coefficient to balance the emphasis on risk estimation and gradation. Experiments show that at $\lambda=0.3$ the model performance is optimal with close and high scores for both gradation and survival risk estimation together for both cohorts (supplementary Figure 8 and supplementary Table 1). Thus for all results reported in the paper, we use $\lambda=0.3$. For uncertainty estimation using conformal prediction, we choose the error rate $\alpha=0.1$. Further details of model architecture, loss function and training methodology are provided in Methods.
Figure 3: Explainability analysis on TCGA-GBMLGG data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomic data for chosen regions marked with different coloured rectangles.
2.2 Similarity evaluation of synthesized and real transcriptomic data
The transcriptomic data synthesized by our model, PathoGen, is compared to the real patient transcriptomes to assess similarity. The Spearman correlation between real and synthesized transcriptomic features comprising all genes evaluates to 0.71 for TCGA-GBMLGG ($p-value=1.18\times 10^{-67}$) and 0.72 ($p-value=3.65\times 10^{-69}$) for TCGA-KIRC, and the mean absolute error (MAE) evaluates to 0.1 for both cohorts (Figures 2.a.i, 2.b.i and supplementary Table 2). The gene expression levels for the gene groups show no significant differences between real and synthesized transcriptomes, for both cohorts (unpaired $t$-test $p$-value $>0.05$, supplementary Table 2). Though there exists a high correlation and low MAE between synthetic and real transcriptomic data, there is some variation across the gene sets (Figures 2.a.i, 2.b.i and supplementary Table 2). Notably, synthesized and real oncogene expression levels are highly correlated across both cohorts (Figures 2.a.i, 2.b.i). This is particularly beneficial for cancer grading or survival estimation as oncogenes when mutated or overexpressed serve as biomarkers that drive tumour cell proliferation and help in predicting cancer prognosis Shay [1999]. Thus we conclude that the synthesized transcriptomic data are highly correlated with the real transcriptomic data and we gain confidence towards using the former in the predictive pipeline.
2.3 Performance evaluation of synthesized transcriptomic data in multimodal prediction
2.3.1 Adding synthesized transcriptomic features with digital histopathology improves prediction
We next analyse whether PathoGen synthesised transcriptomes significantly add value to automated gradation and survival risk estimation, over and above whole slide images (WSIs) alone. Evaluation metrics used for grade prediction and survival risk estimation are the Area Under Curve (AUC) and the concordance index (C Index) respectively. Using WSIs alone we obtain an AUC of 0.823 (gradation) and C Index of 0.842 (survival risk) for TCGA-GBMLGG and an AUC of 0.714 (gradation) and a C Index of 0.671 (survival risk) for TCGA-KIRC (Figures 2.a.ii, 2.b.ii). The performance significantly improves when PathoGen synthesized transcriptomics are used alongside WSIs (Wilcoxon rank-sum test $p$-value $<0.05$), resulting in an AUC of 0.890 (gradation) and C Index of 0.861 (survival risk) for TCGA-GBMLGG and an AUC of 0.773 (gradation) and C Index of 0.681 (survival risk) for TCGA-KIRC (Figures 2.a.ii, 2.b.ii). Thus, in the absence of real transcriptomic data, adding synthesized transcriptomic data significantly improves diagnostic and prognostic predictions compared to using only WSI features.
2.3.2 Comparison of predictive performance of synthesized and real transcriptomic data
We next compare gradation and survival risk predictions based on WSIs and PathoGen synthesized transcriptomic data to predictions based on WSIs and real transcriptomic data. Combining real transcriptomic data with WSIs yields an AUC of 0.907 (gradation) and C Index of 0.866 (survival risk) for TCGA-GBMLGG and an AUC of 0.778 (gradation) and C Index of 0.697 (survival) for TCGA-KIRC (Figures 2.a.ii, 2.b.ii) and the corresponding result using synthesized transcriptomic data is provided in Section 2.3.1. Remarkably, model performance using real transcriptomic data is not significantly different to using PathoGen synthesized transcriptomes (Wilcoxon rank-sum test $p$-value $>$ 0.05) for both gradation and survival risk estimation, in both TCGA cohorts (supplementary Tables 4, 5). Thus, the transcriptomic data synthesized by our model, PathoGen, performs comparably for gradation and survival risk estimation with no significant difference on average from real transcriptomic data.
2.3.3 Comparison with state-of-the-art methodologies
We also compare the performance of our methodology with the state-of-the-art models for using real transcriptomic data and WSI patches. For TCGA-GBMLGG, the C Index evaluated for GSCNN Mobadersany et al. [2018b] is 0.781, DeepAttnMISL Yao et al. [2020] is 0.734, MCAT Chen et al. [2021] is 0.817, Pathomic fusion Chen et al. [2020] is 0.826 whereas our methodology yields a C Index of 0.866 using real transcriptomic data and a C Index of 0.861 using synthesized transcriptomic data for survival risk estimation. For gradation, Pathomic fusion Chen et al. [2020] has an AUC of 0.906 whereas our methodology has an AUC of 0.907 on real transcriptomic data and an AUC of 0.890 on synthesized transcriptomic data. For TCGA-KIRC, the C Index evaluated for Pathomic fusion Chen et al. [2020] is 0.720 whereas for our methodology is 0.697 on real transcriptomic data and 0.681 on synthesized transcriptomic data. However, Pathomic fusion uses ROIs annotated by expert pathologists for predicting grades, while we use transcriptomic features to co-attend to the WSI patch embeddings without requiring any expert intervention, thus reducing expert annotation overheads. We could only compare our model with Pathomic fusion for TCGA-GBMLGG gradation and TCGA-KIRC survival risk estimation as other models did not report the respective performances. Since our model performs better or at least comparably with the state-of-the-art models using real transcriptomic data, and we observe that the predictive performance between using real or synthesized transcriptomic data is very close with no significant difference for the test set population (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5), we may conclude that we achieve state-of-the-art performance using gene expression levels synthesized by our model, PathoGen.
Figure 4: Explainability analysis on TCGA-KIRC data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomic data for chosen regions marked with different coloured rectangles.
2.4 Transparency evaluation for clinical interpretability
2.4.1 Visualization of intra-tumour heterogeneity using synthesized transcriptomic data
To visualize intra-tumour heterogeneity, we use our trained model to predict grade and survival risk for individual WSI patches independently, in a distributed manner, allowing identification of regions of varied grade/survival risk importance within a single WSI (Figures 3.b, 4.b and supplementary Figures 10.b, 11.b) and survival estimates (Figures 3.c, 4.c and supplementary Figures 10.c, 11.c). Illustrative images are shown for patients of ground truth grade III, survival time 10.71 months, survival status alive (Figure 3); ground truth grade IV, survival time 26.08 months, survival status deceased (supplementary Figure 10) belonging to TCGA-GBMLGG, and for patients of ground truth grade I, survival time 33.99 months, survival status deceased (Figure 4); ground truth grade II, survival time 22.78 months, survival status alive (supplementary Figure 11) belonging to TCGA-KIRC. Some regions of interest are highlighted by bounding boxes of colours green, cyan and blue. It is evident that higher-grade regions are correlated to higher risks, as is expected (Figure 4.e). We find that considering WSIs at higher resolution helps to obtain more precise predictions for tissue regions, by obtaining distributed predictions at different magnification levels corresponding to a WSI, as is expected (supplementary Figure 9).
Further, we consider the aggregate of the heterogenous (distributed) predictions and compare it with the grade and survival risk predicted by the model considering all WSI patches as input (non-distributed predictions). For TCGA-GBMLGG, the distributed performance is not significantly different to a one-time prediction with all WSI patches (ANOSIM test jointly over gradation and survival risk estimation $p$-value $>$ 0.05, Figure 2.a.iii). However, for TCGA-KIRC, performing distributed prediction improves the performance significantly (ANOSIM test jointly over gradation and survival risk estimation $p$-value $<$ 0.05, Figure 2.b.iii). Thus we see from the results that there is indeed a distributed variation of grade and risk when predicted locally for patches and mapped onto the WSIs as a heatmap, both within and across cohorts, which may provide a visual insight to the human expert regarding cancer heterogeneity.
2.4.2 Visualisation of how transcriptomic data map to histopathology images
To add explainability, we compare the co-attention maps between transcriptomic data and the corresponding WSI patches, obtained using real and synthesized gene expression levels. TCGA-GBMLGG obtains a Spearman correlation of 0.939 and MAE of 2.07 and TCGA-KIRC obtains a Speraman correlation of 0.851 and MAE of 0.59 (Figures 2.a.iv, 2.b.iv and supplementary Table 3). A higher correlation with a high corresponding MAE would denote that the range of co-attention values is large for the particular gene group but the co-attention map is still correlated. For both cohorts, co-attention maps attributing to oncogenes are most correlated followed by tumour suppressor genes and transcription factor genes (Figures 2.a.iv, 2.b.iv and supplementary Table 3). Oncogenes drive cell division and survival, thus promoting cancer, inactivation of tumour suppressor genes can lead to cancer by failing to regulate uncontrolled cell growth, and transcription factor genes influence cell fate and cancer progression by regulating gene expressions, including those of tumour suppressor and oncogenes Hanahan and Weinberg [2000]. As these gene types happen to be strongly indicative of cancer diagnosis and prognosis, having co-attention maps of real and synthesized transcriptomic data to be most correlated for these groups explains the reason why the predicted grade and survival estimates using real and synthesized gene expression levels are also close.
Further, we study the contribution of synthesized transcriptomic features in co-attending to the WSI patch embeddings. We observe that there are variances in the percentage contribution of individual gene groups in co-attention across both cohorts (Figures 2.a.v,2.b.v). For both cohorts, the transcription factor genes and oncogenes have the highest co-attention values with the WSI patches for both real and synthesized gene expression levels and hence contribute the most to the prediction (Figures 3.d, 2.a.v, 4.d, 2.b.v and supplementary Figure 10.d, 11.d). Activation of oncogenes is known to drive tumour proliferation Shay [1999], hence having these as one of the higher contributors to co-attention would be aligned with clinical expectations. Also since transcription factors regulate gene expression levels to influence cancer progression, having these genes as another primary contributor in co-attention explains the reason behind our high prediction accuracy.
We zoom on specific regions of the WSI and compare the distributed grades, distributed risk and the co-attention maps between synthesized transcriptomic features and the WSI patches (Figures 3.e, 4.e and supplementary Figure 10.e, 11.e). The chosen regions are highlighted by bounding boxes of colours green, cyan and blue. We observe that true grade regions correspond to intermediate to high attention scores of transcription factors, and lower-grade regions are highly related to low attention values, as evident from the regions highlighted by the blue bounding box (Figures 3.c, 4.c) Since, transcription factors regulate the gene expression levels for tumour proliferation Hanahan and Weinberg [2000], such correlation between co-attention value and predicted grade explains the reason behind accurate predictions.
Figure 5: Uncertainty quantification on the TCGA-GBMLGG dataset. (a) Performance of survival risk estimation and gradation for various demographic groups using real and synthesized transcriptomic data. (i) comparison between prediction using real and synthesized data (ii) performance for age groups (iii) performance for gender categories (iv) performance for survival status (censorship) categories (b) Plot of overall prediction results for fairness evaluation.
2.5 Uncertainty quantification for clinical reliability
2.5.1 Quantifying uncertainty
We observe that the uncertainty profile of using synthesized transcriptomic data matches with that of using real transcriptomic data for gradation and survival risk estimation, for different patient demographics and stratification across both cohorts, by performing conformal prediction (CP) as described in Methods. For the entire test set population in TCGA-GBMLGG cohort, gradation uncertainty increases from 0.407 to 0.413 when synthesized data is used and the survival risk estimation uncertainty reduces from 0.446 to 0.438 (Figure 5.a.i). For TCGA-KIRC, gradation uncertainty increases from 0.292 to 0.334 and the survival risk uncertainty increases from 0.898 to 0.909 (Figure 6.a.i). However, the changes in uncertainty are not significant proving the similarity in uncertainty profile (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). Similarly for the different demographic groups for both cohorts, the uncertainty profile remains similar for real and synthesized transcriptomic data with no significant difference in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). It is observed that for a few population groups across both cohorts, the required coverage has not been met, because of less number of samples belonging to the calibration set. But for many such cases, using synthesized data has increased the coverage at the cost of an insignificant increase in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, Figures 6.a.i, 6.a.ii). As risk estimation directly uses the transcriptomic data for prediction and not just for co-attention, such differences in obtained coverage are expected even for minimal differences in the gene expression levels between the synthesized and real transcriptomic features. Thus, the uncertainty estimates help to understand the extent to which we may rely on the predictions, like the survival risk bounds. Additionally, the cancer grade conformal sets provide a range of alternative grades to consider before making the final decision.
2.5.2 Fairness of predictions
Next, we investigate algorithmic fairness, by assessing whether the uncertainty in gradation and survival risk predictions shows demographic biases. We observe that the predicted risk is less for patients who are reported to be alive and more for deceased patients, which is as expected (Figures 5.b, 6.b). Also, the uncertainty of survival risk estimation decreases with increasing risk, which is expected because the actual survival time of patients who are alive is beyond the reported survival time. Thus the model estimated survival risk should correspond to a time beyond the reported survival time and have a higher uncertainty measure when compared in terms of reported survival time. It may also be observed that the model tends to correlate higher risks with the highest grade, grade IV, for both cohorts. However, the trend of gradation uncertainty is different for the two cohorts. For TCGA-GBMLGG, the uncertainty of grade prediction is more for grade III, least for grade IV, and intermediate for grade II (supplementary Table 4 and Figure 5.b). But for TCGA-KIRC, the uncertainty is less for grade I and grade II with low estimated survival risks but high for all other grades (supplementary Table 4 and Figure 6.b). In general, the uncertainty of survival risk estimation is always greater if the survival time reported for a living patient is less (Figures 5.b, 6.b). Also, the uncertainty of prediction for alive patients is always high, except for gradation for TCGA-KIRC (Figures 5.a.iv, 6.a.iv). The uncertainty of grade prediction is more for male patients than females for both cohorts, whereas, for survival risk estimation, the uncertainty is more for females in TCGA-GBMLGG and males in TCGA-KIRC (Figures 5.a.iii, 6.a.iii). The uncertainty of risk estimation decreases with age for both cohorts (Figures 5.a.ii, 6.a.ii). For TCGA-GBMLGG, the uncertainty of grade prediction increases with decreasing magnification but remains similar for all magnification levels for TCGA-KIRC where magnification level 1 corresponds to the highest magnification and level 3 the lowest (supplementary Table 4). Thus these results provide an analysis from a health equity and predictive fairness point of view as to how the performance varies across patient demographics.
Figure 6: Uncertainty quantification on the TCGA-KIRC dataset. (a) Performance of survival risk estimation and gradation for various demographic groups using real and synthesized transcriptomic data. (i) comparison between prediction using real and synthesized data (ii) performance for age groups (iii) performance for gender categories (iv) performance for survival status (censorship) categories (b) Plot of overall prediction results for fairness evaluation.
3 Methods
The paper aims to use synthesized transcriptomic data for the automated selection of whole slide image (WSI) features for risk estimation and gradation of cancer cases. The proposed methodology is illustrated in Figure 1 of the main paper. The WSI is patchified and the corresponding patch embedding is obtained using UNI Chen et al. [2024]. Our novel model, PathoGen, generates transcriptomic features from the patch embeddings. Together the patch embeddings and the gene embeddings obtained from the transcriptomic features are thus used by MCAT_GR for gradation and risk estimation. In the following subsections, we discuss the steps of the proposed methodology in detail.
3.1 Preprocessing
Histopathology data. The diagnostic WSIs for a patient are divided into 224 $\times$ 224 patches, as required for obtaining patch embeddings using the pre-trained foundation model, UNI. The magnification level of the WSI is randomly chosen from the available options to make the model magnification level agnostic. The patches whose mean intensity normalized between 0 and 1 is more than 0.8 are excluded. It is observed that such patches consist mostly of the background region and hence, they do not contribute to the prediction. The 224 $\times$ 224 patches are provided as input to UNI to obtain the patch embeddings of dimension 1 $\times$ 1024. UNI Chen et al. [2024] is a pre-trained vision encoder for histopathology developed using private datasets, and thus is not based on the public datasets used in our experiments.
Transcriptomic data. As proposed in MCAT Chen et al. [2021], the corresponding transcriptomic data of the patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors. The gene expression levels are z-score normalized if already not done. Our model, PathoGen, is trained to synthesize such transcriptomic data from the corresponding WSI. Embeddings of dimension 1 $\times$ 1024 are obtained for the gene expression levels of each category using the gene encoders.
Figure 7: Model Architecture. (a) Architecture of our novel diffusion model, PathoGen, used for synthesizing gene expression levels from histopathology images. (b) Architecture of the modified MCAT model for gradation and risk estimation, MCAT_GR. (c) Detailed architecture of PathoGen transformer. (d) Architecture of gene encoder that converts gene expression level to gene embeddings. (e) Representation of genomic-guided co-attention function. (f) Architecture of gene decoder that converts gene embeddings back to gene expression levels.
3.2 Synthesizing transcriptomic data from histopathology images
Background of diffusion
The process of diffusion is represented as a Markov chain where in the forward process noise is added in each step to degrade the the data for a predetermined timestep $T$ Sohl-Dickstein et al. [2015]. Diffusion models are trained to predict the noise introduced in each step from a noisy sample. In the reverse process, the noise is given as input to the diffusion model and denoising is performed for $T$ timesteps to obtain the synthesized data Ho et al. [2020].
The forward process begins with a data sample $\mathbf{x}_{0}$ and gradually adds Gaussian noise at each timestep $t$ Song et al. [2020] following the equation,
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\beta_{t}$ is a predefined noise variance schedule. For any given timestep $t$, the sample $\mathbf{x}_{t}$ is obtained directly from the initial data $\mathbf{x}_{0}$ as per the following equation,
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (2)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ Ho et al. [2020]. This simplifies to the equation,
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (3)
with $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
The reverse process aims to denoise the sample back to the ground truth data. A learnable model is used to predict the noise at each step to obtain the sample $\mathbf{x}_{t-1}$ from $\mathbf{x}_{t}$ Song et al. [2020].
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\sigma_{\theta}(\mathbf{x}_{t},t)),$ (4)
where $\mu_{\theta}$ and $\sigma_{\theta}$ are learned parameters and given the following equations,
$\mu_{\theta}(\mathbf{x}_{t},t):=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),$ (5)
$\sigma_{\theta}(\mathbf{x}_{t},t):=\tilde{\beta}_{t}^{\frac{1}{2}},\quad t\in\{T,T-1,\dots,1\}.$ (6)
The model is trained to minimize the difference between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}$ using the simplified equation,
$L_{diff}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].$ (7)
3.2.1 Architecture of PathoGen
The architecture of our model, PathoGen, to generate transcriptomic features from whole slide images is illustrated in Figure 7.a with further details in Figures 7.c, 7.d, 7.e and 7.f. To obtain the transcriptomic features at a timestep $t-1$, $\mathbf{x}_{t-1}$, the transcriptomic features at timestep $t$ and the WSI path embeddings are provided as input. The process is performed for T timesteps, where $\mathbf{x}_{T}$ is noise and $\mathbf{x}_{0}$ is the synthesized transcriptomic data. The embedding of the gene expression levels in the transcriptomic data at $\mathbf{x}_{t}$ is obtained using the gene encoder consisting of four linear layers with ELU activation as illustrated in Figure 7.d. The gene embedding and the corresponding patch embedding obtained using UNI Chen et al. [2024] goes as input to the PathoGen transformer, illustrated in Figure 7.c. The PathoGen transformer comprises genomic-guided co-attention illustrated in Figure 7.e., proposed in MCAT Chen et al. [2021], and the transformer encoder layers Vaswani et al. [2017]. Genomic-guided co-attention is inspired by the self-attention introduced in transformer Vaswani et al. [2017]. The gene embeddings attend to the WSI embeddings to produce co-attended embeddings which comprises of the most relevant combined features learned while training. Co-attention is performed thrice to ensure that the generated transcriptomic features have correspondence to the input WSI. Further, the transformer embeddings are decoded to gene expression levels using specific gene decoders for the different gene groups. The gene decoders comprise linear layers to ELU activation except the last, as illustrated in Figure 7.f.
3.2.2 Training and inference
Algorithm 1 gives the training algorithm. The model is trained with the loss function in equation 7 for $T=1000$ timesteps with a learning rate of $1\times 10^{-4}$. For training a sample is considered at a random intermediate timestep $t$ sampled uniformly. Algorithm 2 specifies the inference algorithm. The inference is performed for $T=1000$ timesteps by initially providing random noise sampled from mean zero, unit variance Gaussian distribution as input using equation 4.
Algorithm 1 PathoGen Training
repeat
     $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$
     $t\sim\text{Uniform}(\{1,\dots,T\})$
     $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
     Take gradient descent step on $\nabla_{\theta}\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}$
until converged
Algorithm 2 PathoGen Inference
$\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$
for $t=T,\dots,1$ do
     $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $\mathbf{z}=0$
     $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z}$
end for
return $\mathbf{x}_{0}$
3.3 Gradation and risk estimation
The synthesized transcriptomic data using our model, PathoGen, along with the corresponding WSIs is used for gradation and risk estimation, using the model MCAT_GR.
3.3.1 Architecture
The architecture used for gradation and risk estimation using multi-modal histopathology and synthesized transcriptomic data is MCAT Chen et al. [2021] with an additional pipeline for gradation. The modified MCAT architecture, MCAT_GR, is illustrated in Figure 7.b. The features obtained for the pathomic transformer of MCAT_GR are used for gradation by passing through a separate global attention pooling. The transcriptomic feature embeddings from the gene transformer are not used for gradation as we experiment to see that not including the transcriptomic data for grade prediction performs better. The global attention pooling performed for risk estimation and gradation are separate to decrease the dependency between the grade and risk.
3.3.2 Training details
The loss used for training gradation is binary cross entropy loss Mao et al. [2023], $L_{grade}$, and the loss used for training risk estimation is negative log-likelihood survival loss Chen et al. [2021], $L_{risk}$. Thus, the total loss for MCAT_GR is given by the equation,
$L_{MCAT_{G}R}=\lambda\times L_{grade}+(1-\lambda)\times L_{risk}$ (8)
where $\lambda$ is a hyperparameter to control the contribution of gradation and survival loss in the training. The learning rate used is $2\times 10^{-4}$.
3.4 Conformal Prediction and uncertainty estimation
Conformal for gradation
Conformal prediction is performed for gradation to provide the clinicians with a calibration set of the most probable grade options. Marginal conformal prediction ensures that the coverage of prediction obtained for a population is more than $1-\alpha$, where $\alpha$ is a chosen error rate Angelopoulos and Bates [2021]. The conformal condition for gradation is given by the equation,
$1-\alpha\leq\text{I\kern-1.49994ptP}{\,(Y_{test}\in C(X_{test}))}$ (9)
where $C$ represents the conformal set consisting of probable prediction classes to gain the desired coverage of at least $1-\alpha$. Algorithm 3 specifies the procedure for performing conformal gradation Angelopoulos and Bates [2021]. The uncertainty for gradation is computed using the conformal set by the equation,
$uncertainty=\frac{|C|}{N}\times\frac{\Delta_{max}}{N-1}$ (10)
where $|C|$ represents the cardinality of the conformal set, $N$ represents the number of classes and $\Delta_{max}$ stands for the maximum difference between the classes in the conformal set.
Algorithm 3 Conformal prediction for gradation
Input: Calibration samples $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$, error rate $\alpha$triangleright $y_{i}$ is the true grade
Step 1: Compute conformal scores:
for each calibration sample $i$ do
     $\text{score}_{i}\leftarrow 1-p_{i_{y_{i}}}$triangleright where $p_{i_{y_{i}}}$ is the predicted probability of the true class $y_{i}$
end for
Step 2: Compute adjusted quantile:
$q\leftarrow\frac{\lceil(n+1)\cdot(1-\alpha)\rceil}{n}$
$\hat{q}\leftarrow\text{quantile}(\{\text{score}_{i}\},q)$
Step 3: Form calibration sets:
for each test sample do
     if $p_{j}\geq 1-\hat{q}$ then
         Include class $j$ in the calibration set $\triangleright$ where $p_{j}$ is the predicted probability of the sample belonging to class $j$
     end if
end for
Algorithm 4 Conformal prediction for risk estimation
Input: Calibration samples $\{\mathbf{x}_{i},r_{i}\}_{i=1}^{n}$, error rate $\alpha$
$\triangleright$ $r_{i}$ is the upper bound risk of the true time bin representing maximum risk
Step 1: Compute conformal scores:
for each calibration sample $i$ do
     $\text{score}_{i}\leftarrow|r_{i}-risk_{est_{i}}|$triangleright where $risk_{est_{i}}$ is the predicted risk
end for
Step 2: Compute adjusted quantile:
$q\leftarrow\frac{\lceil(n+1)\cdot(1-\alpha)\rceil}{n}$
$\hat{q}\leftarrow\text{quantile}(\{\text{score}_{i}\},q)$
Step 3: Form calibration sets:
for each test sample do
     $risk_{lb}\leftarrow risk_{pred}-\hat{q}$
     $risk_{ub}\leftarrow risk_{pred}+\hat{q}$
     if $risk_{lb}\leq t_{lb}$ or $t_{ub}\leq risk_{ub}$ then
         Include time bin $t$ in the calibration set
$\triangleright$ where $t_{lb}$ and $t_{ub}$ are the lower bound and upper bound risks corresponding to time bin $t$
     end if
end for
3.4.1 Conformal for risk estimation
Conformal prediction is performed to estimate the upper and lower bounds of the predicted risk. The marginal conformal condition for risk estimation is given by the equation,
$1-\alpha\leq\text{I\kern-1.49994ptP}{\,(risk_{lb}\leq risk_{est}\leq risk_{ub})}$ (11)
where $risk_{est}$ is the estimated risk, and $risk_{lb}$ and $risk_{ub}$ are the lower and upper bound of risks obtained from conformal prediction. For risk estimation, the survival time in months is divided into 4 time bins. The time bins covered between the lower and upper bound of the risk give the conformal set $C$. Algorithm 4 mentions the detailed procedure for conformal risk estimation Angelopoulos and Bates [2021]. The uncertainty for risk estimation is further computed as,
$uncertainty=\frac{|C|}{N}\times\frac{|risk_{ub}-risk_{lb}|}{(-1)-(-5)}$ (12)
where N=4, the number of time bins, -1 is the highest possible risk and -5 is the lowest possible risk score.
4 Discussion
This work has demonstrated, through extensive experiments and state-of-the-art results on two publicly available cancer datasets from TCGA, that using generative AI to synthesize a set of transcriptomic signatures helps to improve multimodal prediction of cancer gradation and patient survival risk when combined judiciously with features learned with a vision transformer based foundation model from digital pathology images. However, in addition to the strength of innovation and the strong performance, in order to increase the trustworthiness of our PathoGen model for clinicians and pathologists, we have presented detailed supporting evidence from the point of view of transparency (mapping back the gradation and risk estimates in a distributed manner onto the WSI images for visual interpretation), reliability (uncertainty quantification by providing conformal sets for gradation and conformal bounds for risk estimation) and fairness (analysis of predictive variation across patient demographics with respect to sex, age, etc). For reproducibility of research and ethos of open science, we have made the code fully available online through our GitHub repository at https://github.com/Samiran-Dey/PathoGen.
Remarkably, we observe that the predictive performance of a multimodal model, combining PathoGen synthesized transcriptomic markers with digital pathology, is comparable to using real transcriptomic signatures. This does not replace the need for collection of transcriptomic data, since the predictive model only focuses on few pan-cancer general signatures, rather than the full transcriptome. However, this does open the door towards an exciting future line of work, that is to investigate further on whether this approach can help in estimating testing the added value of collecting transcriptomic data for the diagnostic/prognostic assessment of each patient when only routine H&\&&E stained slides are available. For future work, we envision that PathoGen can be studied as a screening tool, to guide the targeted collection of expensive transcriptomic data modalities, by potentially selecting patients with highest correlation with the predictive output of the model. This would of course need a dedicated follow up study to validate which would include patient response to treatment when selected in this manner.
Acknowledgment
S Dey is funded by the IIT KGP AI4ICPS Chanakya Fellowship. We acknowledge the Indo-Swedish DBT-Vinnova project, BT/PR41025/Swdn/135/9/2020, for supporting this research. T Chakraborti is supported by the Turing-Roche Strategic Partnership. We are particularly grateful to Prof Brendanc McCane from the Alan Turing Institute and Mr. Chris Harbron from Roche Pharmaceuticals for their insightful comments and constructive feedback throughout this project.
Supplementary Material
Table 1: Table of results for evaluation of the use of synthesized transcriptomic data for grade and survival prediction in four different setting 1) We do not use transcriptomic data at all and attend to the WSI patches randomly; 2) We use the gene expression levels only to co-attend to the WSI patch embeddings; 3) We use the transcriptomic embeddings collectively with co-attended WSI features for survival risk estimation only; 4) We use transcriptomic features for co-attending to WSI patches, and then use the transcriptomic embeddings along with the co-attended features for performing both survival risk estimation and gradation. Setting 3 gives the best mean AUC and C Index for gradation and survival risk estimation across both data cohorts. Further, ANOSIM test for combined risk estimation and gradation between the two cases suggested that the difference in performance between all the cases is significant (p-value $<0.05$). Thus, the architecture of our model is aligned to setting 3. ↑ indicates the larger the better. Best mean scores are marked in bold.
Setting 1 Setting 2 Setting 3 Setting 4
random (not used) only co-attention co-attention, risk co-attention, risk, grade
Risk Grade Risk Grade Risk Grade Risk Grade
$\lambda$ C Index ↑ AUC ↑ C Index ↑ AUC ↑ C Index ↑ AUC ↑ C Index ↑ AUC ↑
TCGA - GBMLGG
0.1 0.835 0.829 0.843 0.878 0.863 0.884 0.848 0.868
0.2 0.832 0.863 0.872 0.869 0.852 0.895 0.864 0.809
0.3 0.842 0.823 0.849 0.886 0.861 0.890 0.859 0.849
0.4 0.844 0.828 0.857 0.885 0.864 0.886 0.858 0.829
0.5 0.856 0.847 0.861 0.890 0.846 0.864 0.830 0.844
0.6 0.847 0.873 0.865 0.877 0.847 0.883 0.820 0.733
0.7 0.861 0.836 0.865 0.867 0.855 0.888 0.879 0.843
0.8 0.862 0.745 0.857 0.856 0.836 0.811 0.808 0.731
0.9 0.850 0.810 0.852 0.801 0.845 0.809 0.801 0.724
mean 0.843 0.820 0.848 0.839 0.854 0.847 0.825 0.796
TCGA - KIRC
0.1 0.626 0.729 0.666 0.713 0.703 0.773 0.576 0.741
0.2 0.709 0.713 0.658 0.768 0.650 0.778 0.516 0.735
0.3 0.671 0.714 0.688 0.763 0.681 0.773 0.548 0.746
0.4 0.701 0.712 0.701 0.770 0.675 0.772 0.513 0.740
0.5 0.662 0.770 0.605 0.766 0.658 0.769 0.509 0.747
0.6 0.672 0.763 0.654 0.767 0.667 0.763 0.516 0.743
0.7 0.677 0.739 0.673 0.768 0.643 0.760 0.512 0.750
0.8 0.708 0.763 0.650 0.774 0.613 0.765 0.551 0.753
0.9 0.607 0.754 0.592 0.760 0.575 0.757 0.573 0.742
mean 0.617 0.742 0.629 0.737 0.639 0.765 0.575 0.742
Figure 8: Ablation for $\lambda$. $\lambda=0$ and $\lambda=1$ correspond to the extreme cases of equation $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$ where the model is trained only using gradation loss and only using survival loss respectively. Using only gradation loss leads to poor performance of risk estimation when $\lambda=0$ and using only risk estimation loss leads to a poor gradation performance. Learning both grade and survival risk together helps the model to correlate between the two and learn better. This is expected as one of the losses essentially becomes unused, so we decide to ignore the extreme cases of $\lambda=0$ and $\lambda=1$. Also, we observe that for $\lambda=0.3$ the model reaches an optimal performance where both AUC and C Index scores are close and high together for both cohorts.
Table 2: Table of results for comparison of synthesized and real transcriptomic features. The correlation coefficient is calculated using Spearman correlation. The p-values are evaluated by performing an unpaired T-test for the aggregate gene expression level of each group. ↑ indicates the larger the better and ↓ indicates the smaller the better. A higher correlation with a high corresponding MAE would denote that the range of gene expression levels is large for the particular group.
Dataset TCGA - GBMLGG TCGA - KIRC
Gene type MAE ↓ Correlation ↑ p-value ↓ MAE ↓ Correlation ↑ p-value ↓
Tumor Suppressor Genes 0.137 0.472 4.32e-06 0.116 0.629 9.21e-69
Oncogenes 0.085 0.770 8.05e-53 0.074 0.817 7.08e-54
Protein Kinases 0.084 0.749 4.83e-79 0.078 0.785 6.48e-84
Cell Differentiation Markers 0.069 0.724 8.55e-50 0.070 0.777 1.38e-53
Transcription Factors 0.122 0.615 4.52e-47 0.117 0.666 6.107e-52
Cytokines and Growth Factors 0.062 0.771 2.43e-66 0.095 0.610 2.48e-72
All genes 0.098 0.713 1.18e-67 0.097 0.717 3.65e-69
Figure 9: Predicted intra-tumour heterogeneity at different magnification levels (a) A whole slide image (WSI) of ground truth grade II, survival time 62.62 months and alive survival status belonging to a male patient of 47 years. (b) A WSI of ground truth grade IV, survival time 54.61 months and dead survival status belonging to a 55 years old female patient. The WSIs belong to TCGA-KIRC. It is observed that for higher resolution images the distributed predictions are more accurate.
Figure 10: Explainability analysis on TCGA-GBMLGG data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomes and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
Figure 11: Explainability analysis on TCGA-KIRC data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomes and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
Table 3: Table of results for co-attention maps generated while using synthesized and real transcriptomic features. The correlation coefficient is calculated using Spearman correlation. ↑ indicates the larger the better and ↓ indicates the smaller the better. A higher correlation with a high corresponding MAE would denote that the range of gene expression levels is large for the particular group.
Dataset TCGA - GBMLGG TCGA - KIRC
Gene type MAE ↓ Correlation ↑ MAE ↓ Correlation ↑
Tumor Suppressor Genes 0.801 0.998 0.087 0.946
Oncogenes 2.302 0.999 0.460 0.995
Protein Kinases 2.526 0.894 0.429 0.881
Cell Differentiation Markers 2.305 0.969 0.319 0.610
Transcription Factors 1.973 0.997 1.865 0.990
Cytokines and Growth Factors 2.505 0.531 0.387 0.404
All genes 2.069 0.939 0.591 0.851
Table 4: Table of results across various groups for grade prediction. p-values are computed using the Wilcoxon rank-sum test to compare the scores for grades predicted using real and synthesized transcriptomic data. ↑ indicates the larger the better and ↓ indicates the smaller the better. Equivalent, better or close ($<0.01$) performance with synthesized data is marked in bold.
Real Synthesized p-value ↑
Category AUC ↑ uncertainty ↓ coverage ↑ AUC ↑ uncertainty ↓ coverage ↑ grade uncertainty
Dataset: TCGA - GBMLGG
overall 0.907 0.407 0.981 0.890 0.413 0.981 0.995 0.935
male 0.915 0.439 0.985 0.896 0.444 0.985 0.975 0.920
female 0.896 0.184 0.921 0.886 0.202 0.921 1.000 0.921
age $<$ 40 years 0.883 1.000 1.000 0.849 1.000 1.000 0.736 1.000
age from 40 to 60 years 0.928 0.459 1.000 0.921 0.450 1.000 0.804 0.996
age $>$ 60 years 0.909 0.722 0.958 0.906 0.708 0.958 1.000 0.934
magnification 1 0.902 0.195 0.966 0.885 0.172 0.931 1.000 0.652
magnification 2 0.952 0.754 1.000 0.936 0.754 1.000 0.811 1.000
magnification 3 0.852 0.901 1.000 0.836 0.901 1.000 0.795 1.000
alive 0.865 0.990 1.000 0.832 1.000 1.000 0.894 0.881
dead 0.951 0.234 0.973 0.951 0.234 0.973 0.867 1.000
grade II 0.914 0.547 0.974 0.891 0.598 0.974 0.704 0.569
grade III 0.755 1.000 1.000 0.728 1.000 1.000 0.866 1.000
grade IV 0.995 0.034 0.966 0.993 0.034 0.966 0.822 1.000
time bin 1 0.898 0.431 0.976 0.882 0.407 0.963 0.984 0.790
time bin 2 0.786 0.538 1.000 0.784 0.513 1.000 1.000 0.817
time bin 3 1.000 0.037 1.000 1.000 0.074 1.000 1.000 0.691
Dataset: TCGA - KIRC
overall 0.778 0.292 0.922 0.773 0.334 0.948 0.850 0.068
male 0.726 0.303 0.909 0.721 0.337 0.929 0.905 0.246
female 0.864 0.278 0.963 0.864 0.336 0.981 0.878 0.123
age $<$ 40 years 0.963 0.056 1.000 0.852 0.111 1.000 0.513 0.513
age from 40 to 60 years 0.780 0.393 0.976 0.778 0.389 0.964 1.000 0.894
age $>$ 60 years 0.770 0.295 0.879 0.761 0.318 0.894 0.886 0.521
magnification 1 0.784 0.369 0.980 0.780 0.337 0.961 1.000 0.393
magnification 2 0.780 0.317 0.961 0.774 0.376 0.980 0.867 0.125
magnification 3 0.771 0.337 0.941 0.769 0.363 0.941 0.867 0.495
alive 0.826 0.239 0.931 0.822 0.288 0.951 0.905 0.076
dead 0.661 0.402 0.922 0.656 0.428 0.941 0.880 0.495
grade I 0.222 0.500 0.667 0.111 0.500 0.333 1.000 1.000
grade II 0.987 0.016 0.905 0.978 0.000 0.873 0.880 0.878
grade III 0.647 0.283 0.905 0.658 0.333 0.952 0.645 0.164
grade IV 0.504 0.625 0.958 0.499 0.563 0.958 0.711 0.458
time bin 1 0.718 0.366 0.944 0.713 0.384 0.944 0.889 0.565
time bin 2 0.837 0.214 0.883 0.828 0.233 0.900 0.885 0.560
time bin 3 0.848 0.296 1.000 0.835 0.333 1.000 1.000 0.569
time bin 4 0.667 0.111 0.667 0.667 0.167 1.000 1.000 0.513
Table 5: Table of results across various groups for risk estimation. p-values are computed using the Wilcoxon Ranksum Test to compare the scores for grades predicted using real and synthesized transcriptomic data. ↑ indicates the larger the better and ↓ indicates the smaller the better. Equivalent, better or close ($<0.01$) performance with synthesized data is marked in bold.
Real Synthesized p-value ↑
Category C Index ↑ uncertainty ↓ coverage ↑ C Index ↑ uncertainty ↓ coverage ↑ risk uncertainty
Dataset: TCGA - GBMLGG
overall 0.866 0.446 0.933 0.861 0.438 0.923 0.604 0.652
male 0.808 0.375 0.985 0.813 0.376 0.985 0.956 0.956
female 0.795 0.653 0.921 0.795 0.639 0.921 0.350 0.400
age $<$ 40 years 0.524 0.767 0.977 0.540 0.727 1.000 0.119 0.135
age 40 to 60 years 0.848 0.480 0.973 0.845 0.492 0.973 0.775 0.770
age $>$ 60 years 0.738 0.377 1.000 0.742 0.387 1.000 0.741 0.734
magnification 1 0.855 0.735 0.966 0.840 0.725 1.000 0.646 0.810
magnification 2 0.862 0.484 0.974 0.856 0.486 0.974 0.950 0.975
magnification 3 0.873 0.540 0.973 0.858 0.517 0.973 0.578 0.578
alive - 0.368 0.851 - 0.359 0.851 0.237 0.441
dead 0.614 0.304 0.973 0.633 0.314 1.000 0.563 0.563
grade II 0.823 0.582 0.872 0.823 0.544 0.821 0.058 0.054
grade III 0.590 0.655 0.972 0.526 0.659 0.972 0.901 0.866
grade IV 0.509 0.056 0.966 0.539 0.059 0.966 0.489 0.494
time bin 1 0.827 0.252 0.951 0.822 0.259 0.939 0.942 0.893
time bin 2 - 0.804 0.923 - 0.782 0.923 0.209 0.270
time bin 3 - 0.996 1.000 - 0.984 1.000 0.200 0.233
Dataset: TCGA - KIRC
overall 0.697 0.898 0.882 0.681 0.909 0.928 0.300 0.307
male 0.705 0.909 0.909 0.652 0.921 0.960 0.669 0.654
female 0.710 0.855 0.759 0.771 0.865 0.796 0.219 0.311
age $<$ 40 years - 0.955 1.000 - 0.942 1.000 0.275 0.275
age 40 to 60 years 0.612 0.936 0.893 0.591 0.935 0.929 0.266 0.265
age $>$ 60 years 0.672 0.864 0.909 0.681 0.891 0.985 0.781 0.781
magnification 1 0.706 0.932 0.961 0.676 0.947 1.000 0.563 0.588
magnification 2 0.734 0.899 0.902 0.702 0.908 0.961 0.248 0.277
magnification 3 0.694 0.878 0.863 0.676 0.890 0.882 0.408 0.493
alive - 0.934 0.912 - 0.934 0.941 0.083 0.089
dead 0.487 0.845 0.902 0.499 0.874 0.980 0.595 0.597
grade I - 0.976 0.667 - 0.969 1.000 0.275 0.513
grade II 0.835 0.926 0.905 0.841 0.919 0.937 0.139 0.149
grade III 0.612 0.906 0.857 0.573 0.904 0.857 0.550 0.659
grade IV 0.481 0.901 0.958 0.439 0.964 1.000 0.143 0.112
time bin 1 0.690 0.922 0.861 0.714 0.938 0.944 0.441 0.474
time bin 2 0.842 0.864 0.983 0.836 0.881 0.983 0.557 0.502
time bin 3 - 0.389 1.000 - 0.389 1.000 0.411 1.000
time bin 4 - 0.775 1.000 - 0.775 1.000 0.275 1.000